OpenStudy (anonymous):

I need some quick help with an algebra II problem - nothing complicated, just using the interest formula. I do not want the answer - I want help understanding the process. Thank you!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

You decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula \[A(t) = P \left[ \left( 1 + r/n \right)^{1/c} \right]^{cnt}\] where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1 + r/n.

OpenStudy (anonymous):

Here are my givens (sorry, forgot to post them first): You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Price: $39,145 Current interest rate: 3% (0.03)

OpenStudy (anonymous):

@megangray32 @jigglypuff314 @amistre64 @EducateMe @xapproachesinfinity

OpenStudy (anonymous):

Hi @xapproachesinfinity Can you help? :)

OpenStudy (anonymous):

I'll go ahead and post what I have so far.

OpenStudy (anonymous):

A(4) = [39,145(1 + 0.03/12)^12(4)^1/4]^(4)(12)(4) ??

OpenStudy (anonymous):

@xapproachesinfinity Help? :(

OpenStudy (anonymous):

In which case, do I need to set it equal to 1 + r/n? (1 + 0.03/12) = [39,145(1 + 0.03/12)^12(4)^1/4]^(4)(12)(4)

OpenStudy (anonymous):

1.0025 = [39,145(1 + 0.03/12)^1/4]^(4)(12)(4) I had accidentally included ^nt instead of ^1/c. :)

OpenStudy (anonymous):

@jim_thompson5910 @jbreezy622 @Firejay5

OpenStudy (anonymous):

Hey @jbreezy622 ! Can you help?

OpenStudy (anonymous):

@kohai

OpenStudy (anonymous):

I need help with algebra myselflol

OpenStudy (anonymous):

Oh, nooo! I'm hoping I'm on the right track with this. It's the last question on my assignment and I need to make sure I understand the concept before the test. @jbreezy622

OpenStudy (anonymous):

@satellite73 Please help!!

OpenStudy (anonymous):

@mancake1

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@mary.rojas @meg_72 @justuu

OpenStudy (anonymous):

@Krinkov

OpenStudy (anonymous):

@satellite73

OpenStudy (mary.rojas):

sorry,busy extremely :( good luck!

OpenStudy (anonymous):

@Krinkov Can you help with this?

OpenStudy (anonymous):

The "4 times a month" thing is throwing me off. Does that make n = 48, or t = 4, or...???

OpenStudy (anonymous):

@Firejay5 Please help! I think I figured out that n needs to be 52, the number of weeks in a year.